# Likelihood Based Inference for ARMA Models

### v5

@@ 3 Annual Depths of Lake Michigan
In this example, we illustrate how improperly maximized likelihoods can lead to inconsistencies and uncertainty in a real data analysis scenario. Additionally, we show how the common practice of using the estimated standard error for calibrated parameters can misleadingly support the inclusion of model parameters. We consider a dataset containing annual observations on the average depth of Lake Michigan-Huron, recorded the first day of each year from 1860-2014 (Figure 5) (NOAA, 2016). For our analysis, we adhere to the currently accepted standards for data analysis, such as those outlined in Shumway and Stoffer (2017).
Figure 5: Average depth of Lake Michigan-Huron from 1860-2014.
@@
Standard diagnostic tests suggest that it is reasonable to hypothesize the data in Figure 5 could be appropriately modeled as a weakly stationary $\text{ARMA}(p,q)$ process for some non-negative integers $p$ and $q$. While an ARIMA model may also be reasonable for these data, we first consider fitting an ARMA model because we would like to avoid the possibility of over-differencing the data. The next step is to determine appropriate values of $p$ and $q$; after some initial investigation, multiple combinations of $p$ and $q$ seem plausible, and therefore we decide to choose the values of $p$ and $q$ that minimize the AIC. For simplicity, we create a table of AIC values for all possible combinations of $(p,q)\in\{0,1,2,3\}^{2}$ (Table 1). Using the AIC as the model selection criterion, the selected model size is $\text{ARMA}(2,1)$.
@@
Minimizing the AIC (or an alternative information criterion) is not the only accepted approach to order selection. A classical perspective on model selection involves consulting sample autocorrelation plots, partial autocorrelation plots, conducting tests such as Ljung-Box over various lags, studying the polynomial roots of fitted models, and checking properties of the residuals of the fitted models (Box and Jenkins, 1970; Brockwell and Davis, 1991; Shumway and Stoffer, 2017). This approach helps avoid fitting models that are possibly over-parameterized. However, additional computational power and increasing volumes of data have favored automated data analysis strategies that fit many models and evaluate them using a model selection criterion. In principle, a simple model selection criterion such as AIC can address parsimony and guard against over-parameterization as well. Diagnostic inspection can be combined with these automated approaches. For example, a table of AIC values can be generated, and models with promising likelihoods can be explored further (Brockwell and Davis, 1991).
@@ 3.1 Parameter uncertainty
Table 2: Parameter values of $\text{ARMA}(p,q)$ model fit to Lake Michigan-Huron depth data.
@@
However, our simulation studies have suggested that these confidence intervals can be misleading, and that PLCIs are more reliable alternatives. The $95\%$ PLCI for the parameter (Panel (A) of Figure 7) is much larger than the confidence interval created using these standard errors. The steep curve in the immediate vicinity of $\hat{\theta}_{1}$ may explain the small standard error estimates for this parameter and the corresponding tight confidence intervals created using Fisher’s identity matrix. Alternative evidence indicates the potential for nearly canceling roots (Figure 6), in which case the MA(1) term may not be needed in the model.
Figure 6: Inverted AR and MA polynomial roots to the fitted $\text{AR}(1)$ and $\text{ARMA}(2,1)$ models to the Lake Michigan-Huron data using a single parameter initialization.
Both types of confidence intervals considered in this example rely on asymptotic justifications, but we can further investigate the finite sample properties using a simulation study. We fit both $\text{ARMA}(2,1)$ and $\text{AR}(1)$ models to the data, and conduct a boot-strap simulation study by simulating $1000$ datasets from each of the fitted models. We then re-estimate an $\text{ARMA}(2,1)$ model to each of these datasets and record the estimated coefficients. A histogram containing the estimated values of $\hat{\theta}_{1}$ when the data are generated from the $\text{ARMA}(2,1)$ and $\text{AR}(1)$ models fit to the Lake Huron-Michigan data are shown in Panels (B) and (C) of Figure 7, respectively.
The shape of this histogram in Panel (B) mimics that of the profile log-likelihood surface in Panel (A), confirming that a large confidence interval is needed in order to obtain a $95\%$ confidence interval. In Panel (C), a large number of $\hat{\theta}_{1}$ coefficients are estimated near $1$ when the generating model is $\text{AR}(1)$. Combining this result with the nearly canceling roots of the $\text{ARMA}(2,1)$ model (Figure 6), we cannot reject the hypothesis that the data were generated from a AR(1) model, even though the Fisher information standard errors suggest that the data should be modeled with a nonzero $\theta_{1}$ coefficient.
@@
Figure 7: Evidence for an AR(1) model for the Lake Michigan-Huron data. (A) Profile likelihood confidence interval (PLCI) for $\theta_{1}$ which includes the value $\theta_{1}=0$. The vertical dotted line represents the lower end of the approximate confidence interval; all points on the solid black line lie within the confidence interval, and points on the dashed red line are outside the interval. (B) Histogram of re-estimated $\theta_{1}$ values using simulated data simulated from the $\text{ARMA}(2,1)$ model that was calibrated to the Lake Michigan-Huron data. (C) Histogram of re-estimated $\theta_{1}$ values using data simulated from the AR(1) model that was calibrated to the Lake Michigan-Huron data.
@@ 4 Discussion
An important consequence of improved likelihood maximization is better model selection. A common approach to selecting an ARMA model involves fitting different sizes of models and choosing the one that minimizes an information criterion, such as the AIC. Fitting multiple models results in having a higher probability that at least one candidate model was not properly maximized. Since AIC assumes the parameters correspond to maximized likelihoods, enhancements in likelihood maximization can lead to different model selections. Consequently, methodology relying on existing estimation methods—like the popular auto.arima function in the forecast package in R (Hyndman and Khandakar, 2008), which minimizes the AIC of a group of candidate models without explicitly displaying an AIC table—will be impacted by improved estimates.
Parameter estimates corresponding to higher likelihood values are not necessarily better than those with lower likelihood values (Le Cam, 1990). Sometimes, our improved estimates may result in models with nearly canceling roots, parameters near boundary conditions, or otherwise unfavorable statistical properties. On other occasions, our method can rescue a naive optimization attempt from a local maximum having those unfavorable properties. Practitioners should carefully evaluate fitted models to ensure they are appropriate for the data and problem at hand.
@@
Our proposed algorithm for ARMA parameter estimation significantly advances statistical practice by addressing a frequently occurring optimization deficiency. Because existing software can also be leveraged to mitigate the issue, the largest contribution of this work may be highlighting the prevalence of this optimization problem. Traditional random initialization approaches software fail to uniformly cover the entire range of possible models and often produce many initializations outside the accepted range. Our algorithm offers a computationally efficient and practically convenient solution, providing a robust approach to parameter initialization and estimation that ensures comprehensive coverage of possible models. We have shown that it provides a new standard for best practice in the field of time series analysis.
Appendix A Uniform Sampling
A common approach to optimizing a non-convex loss function is to perform the optimization routine with distinct parameter initializations. For ARMA models, picking suitable initialization is a challenging problem that we address with Algorithm 1. An alternative approach would involve sampling each coefficient independently. To see why an independent sampling scheme is not used, consider an AR(2) model. An initialization with parameters $(\phi_{1},\phi_{2})=(1.1,0.1)$ is not a valid initialization—as the polynomial roots lie outside the complex unit circle—whereas $(\phi_{1},\phi_{2})=(1.1,-0.2)$ is perfectly acceptable. Our algorithm accounts for the complex relationship between model parameters when obtaining random initializations.
To visualize why an independent sampling scheme is not used, consider sampling parameters from a Uniform(1, 1) distribution. In Figure 8, we plot inverted roots that are a result of sampling from this distribution for AR(2) and AR(3) models. The figure illustrates that a significant percentage of uniformly sampled parameter initializations lie outside the accepted region, and, critically, the entire region of possible initializations is not well covered by uniform sampling. Picking a uniform distribution with different bounds—or any other independent sampling distribution—results in similar problems.
Figure 8: Inverted roots of 1000 samples of AR(2) and AR(3) coefficients sampled from a U(1, 1) distribution. The red “x”s are points that lie outside the accepted region, which represents $13\%$ of the AR(2) coefficients and $22.6\%$ of the AR(3) coefficients. Increasing the width of the uniform sampling distribution results in a larger fraction outside the accepted region, and decreasing the width results in worse coverage of the range of acceptable parameter values.
Appendix B Python Comparison
Although Algorithm 1 is language agnostic, our implementation is in R. Consequently, all simulation studies for this article use R’s stats::arima function for baseline comparisons. To demonstrate applicability to other software environments, we briefly compare model likelihoods fit using Python’s statsmodels.tsa module against our implementation of Algorithm 1. We generate 100 unique Gaussian ARMA$(2,1)$ models and datasets (each with $n=100$) where R’s stats::arima provides sub-optimal estimates. These observations are used to fit ARMA$(2,1)$ models in Python. Despite these datasets being chosen for sub-optimal results in R, the log-likelihoods in both R and Python are roughly equivalent—a result of both software packages using the same general approach to fitting model parameters.
Our implementation of Algorithm 1 resulted in higher log-likelihoods for $97$ out of the 100 datasets compared to Python. The log-likelihood deficiencies for the remaining three datasets were all smaller than $\epsilon=10^{-5}$, which is smaller than the tolerance level to be considered as an improvement in our other simulation studies. While potentially insignificant, these differences can be eliminated by increasing the number of parameter initializations and the convergence criteria of our algorithm. Directly implementing our algorithm in Python would further eliminate the possibility of these discrepancies entirely.

### v6

@@ 3.2 Annual Depths of Lake Michigan
In this example, we illustrate how improperly maximized likelihoods can lead to inconsistencies and uncertainty in a real data analysis scenario. Additionally, we show how the common practice of using the estimated standard error for calibrated parameters can misleadingly support the inclusion of model parameters. We consider a dataset containing annual observations on the average depth of Lake Michigan-Huron, recorded the first day of each year from 1860-2014 (Fig 5) [25]. The data are provided as part of the arima2 package. We wish to develop an ARMA model for these data, which is a standard task in time series analysis [4].
Fig 5: Average depth of Lake Michigan-Huron from 1860-2014.
@@
Diagnostic tests, such as sample autocorrelation and normal quantile plots for residuals, suggest that it is reasonable to model the data in Fig 5 as a weakly stationary Gaussian $\text{ARMA}(p,q)$ process for some non-negative integers $p$ and $q$. While an ARIMA model may also be reasonable for these data, we first consider fitting an ARMA model because we would like to avoid the possibility of over-differencing the data. The next step is to determine appropriate values of $p$ and $q$; after some initial investigation, multiple combinations of $p$ and $q$ seem plausible, and therefore we decide to choose the values of $p$ and $q$ that minimize the AIC. For simplicity, we create a table of AIC values for all possible combinations of $(p,q)\in\{0,1,2,3\}^{2}$ (Table 2). Using the AIC as the model selection criterion, the selected model size is $\text{ARMA}(2,1)$.
@@
Minimizing the AIC (or an alternative information criterion) is not the only accepted approach to order selection. A classical perspective on model selection involves consulting sample autocorrelation plots, partial autocorrelation plots, conducting tests such as Ljung-Box over various lags, studying the polynomial roots of fitted models, and checking properties of the residuals of the fitted models [2, 22, 4]. This approach helps avoid fitting models that are possibly over-parameterized. However, additional computational power and increasing volumes of data have favored automated data analysis strategies that fit many models and evaluate them using a model selection criterion. In principle, a simple model selection criterion such as AIC can address parsimony and guard against over-parameterization as well. Diagnostic inspection can be combined with these automated approaches. For example, a table of AIC values can be generated, and models with promising likelihoods can be explored further [22].
@@ 3.2.1 Parameter uncertainty
Table 3: Parameter values of $\text{ARMA}(p,q)$ model fit to Lake Michigan-Huron depth data. The same parameters are returned with a single initialization, or when using Algorithms 1 or 2.
@@
However, our simulation studies have suggested that these confidence intervals can be misleading, and that PLCIs are more reliable alternatives. The $95\%$ PLCI for the parameter (Fig 6A) is much larger than the confidence interval created using these standard errors. The steep curve in the immediate vicinity of $\hat{\theta}_{1}$ may explain the small standard error estimates for this parameter and the corresponding tight confidence intervals created using Fisher’s identity matrix. Alternative evidence indicates the potential for nearly canceling roots (Fig 7), in which case the MA(1) term may not be needed in the model.
@@
Fig 6: Evidence for an AR(1) model for the Lake Michigan-Huron data. (A) Profile likelihood confidence interval (PLCI) for $\theta_{1}$ which includes the value $\theta_{1}=0$. The vertical dotted line represents the lower end of the approximate confidence interval; all points on the solid black line lie within the confidence interval, and points on the dashed red line are outside the interval. (B) Histogram of re-estimated $\theta_{1}$ values using simulated data simulated from the $\text{ARMA}(2,1)$ model that was calibrated to the Lake Michigan-Huron data. (C) Histogram of re-estimated $\theta_{1}$ values using data simulated from the AR(1) model that was calibrated to the Lake Michigan-Huron data.
Fig 7: Inverted AR and MA polynomial roots to the fitted $\text{AR}(1)$ and $\text{ARMA}(2,1)$ models to the Lake Michigan-Huron data using a single parameter initialization.
Both types of confidence intervals considered in this example rely on asymptotic justifications, but we can further investigate the finite sample properties using a simulation study. We fit both $\text{ARMA}(2,1)$ and $\text{AR}(1)$ models to the data, and conduct a boot-strap simulation study by simulating $1000$ datasets from each of the fitted models. We then re-estimate an $\text{ARMA}(2,1)$ model to each of these datasets and record the estimated coefficients. A histogram containing the estimated values of $\hat{\theta}_{1}$ when the data are generated from the $\text{ARMA}(2,1)$ and $\text{AR}(1)$ models fit to the Lake Huron-Michigan data are shown in Fig 6B and Fig 6C, respectively.
The shape of this histogram in Fig 6B mimics that of the profile log-likelihood surface in Fig 6A, confirming that a large confidence interval is needed in order to obtain a $95\%$ confidence interval. In Fig 6C, a large number of $\hat{\theta}_{1}$ coefficients are estimated near $1$ when the generating model is $\text{AR}(1)$. Combining this result with the nearly canceling roots of the $\text{ARMA}(2,1)$ model (Fig 7), we cannot reject the hypothesis that the data were generated from a $\text{AR}(1)$ model, even though the Fisher information standard errors suggest that the data should be modeled with a nonzero $\theta_{1}$ coefficient. Diagnostic plots for both $\text{AR}(1)$ and $\text{ARMA}(2,1)$ models are given in Fig 8.
Fig 8: Residual plots for models fit to the Lake Michigan-Huron data. (A) Residuals of fitted AR(1)(1) model. (B) Residuals of fitted ARMA(2,1)(2,1) model.
@@ 4 Discussion
An important consequence of improved likelihood maximization is better model selection. A common approach to selecting an ARMA model involves fitting different sizes of models and choosing the one that minimizes an information criterion, such as the AIC. Fitting multiple models results in having a higher probability that at least one candidate model was not properly maximized. Since AIC assumes the parameters correspond to maximized likelihoods, enhancements in likelihood maximization can lead to different model selections. Consequently, methodology relying on existing estimation methods—like the popular auto.arima function in the forecast package in R [24], which minimizes the AIC of a group of candidate models without explicitly displaying an AIC table—will be impacted by improved estimates.
Our proposed algorithm is supported by existing theory on likelihood evaluation of linear state-space models via the Kalman Filter [16], the same as the current existing standard approach for parameter estimation. The simulation studies that we have conducted, however, demonstrate the importance of considering multiple parameter initializations in order to fully maximize model likelihoods. These simulations provide a conservative estimate of how frequently our algorithm results in improved likelihoods compared to existing standards. A common situation where our algorithm is expected to provide even larger improvements than those reported here is in the presence of missing data, a primary motivator of the likelihood maximization procedure of existing software [14]. In this situation, the well-informed CSS initialization is not available, and the default approach is to initialize at the origin, resulting in a greater need to attempt multiple parameter initializations.
Parameter estimates corresponding to higher likelihood values are not necessarily scientifically preferable to alternative regions of parameter space with lower likelihood values [26]. Sometimes, our improved estimates may result in models with nearly canceling roots, parameters near boundary conditions, or otherwise unfavorable statistical properties. On other occasions, our method can rescue a naive optimization attempt from a local maximum having those unfavorable properties. Alternative forms of parameter estimation may also be preferable in some cases. For example, a series of papers by Chen and Deo show that restricted likelihood estimators have favorable properties compared to maximum likelihood estimators when the AR parameters are near the unit root [8, 9, 10]. Practitioners should carefully evaluate fitted models to ensure they are appropriate for the data and problem at hand.
@@
Our proposed algorithm for ARMA parameter estimation significantly advances statistical practice by addressing a frequently occurring optimization deficiency. Because existing software can also be leveraged to mitigate the issue, the largest contribution of this work may be highlighting the prevalence of this optimization problem. Traditional random initialization approaches software fail to uniformly cover the entire range of possible models and often produce many initializations outside the accepted range. Our algorithm offers a computationally efficient and practically convenient solution, providing a robust approach to parameter initialization and estimation that ensures adequate coverage of all possible models. We have shown that it provides a new standard for best practice in the field of time series analysis.
Supporting information
S1 Appendix
Uniform Sampling. An appendix demonstrating why uniform sampling of ARMA model parameters is impractical.
S2 Appendix
Python Comparison. An appendix comparing our algorithm to existing Python software.
S3 File
Source code All code used to run the simulations studies used in this article and supplement, zipped into a single file.